OpenStudy (camzzzie):

PLEASE HELP WITH DIRECT VARIATION WILL MEDAL AND FAN Determine whether y varies directly with x. If so, find the constant of variation k and write the equation. x y 7 11 8 13 9 15 10 17

OpenStudy (camzzzie):

I know y=kx

OpenStudy (camzzzie):

so k=y/x so I did 11/7=1.7 y=1.7x plugged in 11=1.7(7)+b 11=11.9+b 11-11.9=b b=-.9